OpenStudy (anonymous):

find (dy/dx) when x= 1/4 when y=cos(πx)

OpenStudy (anonymous):

when x = 1/4, y = cos (pi/4) = sqrt(2)/2 dy/dx is the derivative of cos(pi x). The derivative of cos(pi x) using the Chain Rule is - pi sin(pi x) So at x = 1/4, you get - pi sin(pi/4) = - pi (sqrt 2/2) or (- pi sqrt(2)/2)